User name Eric Voutier
Log entry time 00:16:32 on October24,2002
Entry number 88915
New Crew on shift: Alexandre Camsonne, Lisa Kaufman, Eric Voutier
16:00 Start shift with Q1/Left down; controlled access to solve the problem (Ed)
and waiting for the magnet to come back.
16:35 While waiting for Q1_left, the Dipole_left start to ramp down for no clear
reason, after several attempts to confirm and tweak the momentum setting the
dipole finally come back.
16:40 Ready to take beam.
16:48 Start production
16:56 Low gas pressure alarm (Argon and Co2 from the Main control screen),
some of the FPP HV tripped
18:05 Target problem / low pressure D2 alarm / IOC reboot
20:05 Switch off the HV of the pion rejector while trying to swicth on back the
FPP (run 2712), this should affect the last 100 events. FPP is now back in service; start the last run at kinematisc Q3D_d40.
21:00 Kinematics Q3D_d40 ends with run 2715 (efficiency run).
21:10 Ep energy measurement starts.
21:12 Target motion problem
21:49 Target motion solved / Mike Seely came in
22:40 Ep measurement fails, the diagnostic seems to tell that Ep target is broken; need further investigations and checks to conclude.
22:42 Kinematics Q3D_d50 is in place; start taking data on optics target.
23:01 Moving to 15cm LD2
23:10 Back to production; detector analysis indicate
--# Kinematics Q3D_d40
Run # 2709 .......... good run / 1M / some FSD trips / Gas Alarm
Run # 2710 .......... good run / 355 k / target IOC reboot at the end
Run # 2711 .......... good run / 1 M
Run # 2712 .......... good run / 568 k / pion rejector off at the very end
Run # 2713 .......... good run / 1M
Run # 2714 .......... good run / 35 k
Run # 2715 .......... scintillator efficiency run
--# Kinematics Q3D_d50
Run # 2716 .......... pointing measurement
Run # 2717 .......... goes to next shift